# Efficient Nearest Neighbor based Uncertainty Estimation for Natural Language Processing Tasks

### v1

@@ Abstract
Trustworthy prediction in Deep Neural Networks (DNNs), including Pre-trained Language Models (PLMs) is important for safety-critical applications in the real world. However, DNNs often suffer from uncertainty estimation, such as miscalibration. In particular, approaches that require multiple stochastic inference can mitigate this problem, but the expensive cost of inference makes them impractical. In this study, we propose -Nearest Neighbor Uncertainty Estimation (NN-UE), which is an uncertainty estimation method that uses the distances from the neighbors and label-existence ratio of neighbors. Experiments on sentiment analysis, natural language inference, and named entity recognition show that our proposed method outperforms the baselines or recent density-based methods in confidence calibration, selective prediction, and out-of-distribution detection. Moreover, our analyses indicate that introducing dimension reduction or approximate nearest neighbor search inspired by recent NN-LM studies reduces the inference overhead without significantly degrading estimation performance when combined them appropriately.
@@ 1 Introduction
In order to use Deep Neural Networks (DNNs) including Pre-trained Language Models in safety-critical regions, uncertainty estimation is important. By improving the predictive uncertainty, the prediction will be calibrated Guo et al. (2017),111”Calibration” means the confidence of the model aligns with its accuracy. or improve selective prediction performance, which is predictive performance when there is a choice to abstain from model prediction Galil et al. (2023). On the other hand, DNNs often fail to quantify the predictive uncertainty, for example, causing miscalibrated prediction Guo et al. (2017). Such uncertainty estimation performance problems can be mitigated by the PLMs, such as BERT Devlin et al. (2019) or DeBERTa He et al. (2021b), that are self-trained on large amounts of data Ulmer et al. (2022), although, there is still a need for improvement Desai and Durrett (2020).
@@
To solve the problem of uncertainty estimation, multiple probabilistic inferences such as MC Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017) are generally effective. On the other hand, these methods require multiple stochastic inferences for a single data instance, which leads to high computational cost, and makes them impractical for real world application. To obtain reasonable predictive uncertainty without multiple inferences, Temperature Scaling Guo et al. (2017) is generally used, which scales logits with a temperature parameter. Furthermore, density-based methods such as Density Softmax Bui and Liu (2024a) and Density Aware Calibration (DAC) Tomani et al. (2023), which correct the model outputs based on estimated density, have achieved promising very recent years in terms of uncertainty estimation performance and inference cost. However, both Density Softmax and DAC only uses the density of training data. Therefore, we can see that these methods only capture epistemic uncertainty that comes from the knowledge of the model. To improve the uncertainty estimation performance, we also need to consider aleatoric uncertainty that comes from the variance of the data Hüllermeier and Waegeman (2019).
@@
In this study, we propose -Nearest Neighbor Uncertainty Estimation (NN-UE), a new density-based uncertainty estimation method that does not require multiple inferences. NN-UE uses the labels of neighbors obtained from NN search to correct the confidence as illustrated in Figure 1. Our method weights logits according to the distance between the input example and its neighbors in the datastore created by the training data. In addition, the logit weights are further biased by the ratio of the model’s predicted label matched with the labels in neighbors. As a result, our method requires only a single forward inference of the model.
@@
First, our experiments show that NN-UE improves the uncertainty estimation performance of existing baselines in sentiment analysis, natural language inference, and named entity recognition in both in-domain and out-of-domain settings by combining neighbor label information and distances from neighbors. Second, we show that the smaller the number of instances of neighbor search in NN-UE, the better the uncertainty performance tends to be slight. Finally, to solve the latency in NN-UE for token-level tasks, such as sequence-labeling based name entity recognition, we show that approximate nearest neighbor search or dimension reduction in NN-UE improves the inference speed without degrading the uncertainty estimation performance much more, while combining them leads to degrading the uncertainty performance. Our code will be available after acceptance.
@@ 2 Related Work
Uncertainty Estimation
Many prior studies on uncertainty estimation focus primarily on image classification. One approach for improved uncertainty estimation is the sampling-based approach of performing multiple stochastic predictions for one data instance. The approach includes MC Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017). However, their practical application is limited due to the computationally expensive inference. As an alternative to improve uncertainty, the sampling-free approaches do not demand multiple stochastic inferences, e.g., Temperature Scaling Guo et al. (2017), Deep Deterministic Uncertainty Mukhoti et al. (2023), and Posterior Network Charpentier et al. (2020). Very recently, the density-based methods Bui and Liu (2024a); Tomani et al. (2023); Bui and Liu (2024b) show good performance on latency and uncertainty estimation.
@@
-Nearest Neighbor Language Model (NN-LM) Khandelwal et al. (2020) has been proposed, which performs linear interpolation of NN probability based on distance from neighbors and base model probability, in the language modeling task. -Nearest Neighbor Machine Translation (NN-MT) applied the NN-LM framework to machine translation Khandelwal et al. (2021). NN-LM and NN-MT have been successful because they enhance predictive performance through the memorization and use of rich token representations of pre-trained language models and mitigate problems such as a sparsity comes from low-frequency tokens Zhu et al. (2023). The main issue on NN-LM and NN-MT is the inference overhead, and there are several studies to solve this problem. He et al. (2021a) employs datastore compression, adaptive retrieval, and dimension reduction to reduce computational overhead with retaining perplexity. Deguchi et al. (2023) dramatically improves decoding speed by dynamically narrowing down the search area based on the source sentence.
Uncertainty Estimation for Natural Language Processing Tasks
Studies about uncertainty estimation for NLP tasks are limited when compared with those for image datasets. Kotelevskii et al. (2022) has shown excellent performance in classification with rejection tasks and out-of-distribution detection tasks using uncertainty scores using density estimation results. Vazhentsev et al. (2022) performed misclassification detection using Determinantal point processes, spectral normalization, and loss regularization in text classification and NER. Vazhentsev et al. (2023) proposed the new uncertainty method which combined epistemic and aleatoric uncertainty estimation methods for text classification tasks. In this study, we comprehensively investigate the density-based methods such as Density Softmax and DAC in NLP tasks, inspired by the success of memorization and use in NLP region Zheng and Jiang (2022), such as NN-LM and NN-MT. In addition, our study attempts to improve the uncertainty estimation performance by making more effective use of memorized examples, such as introducing label information of neighbor examples.
3 Base Methods
In this section, we describe the density-based methods for uncertainty estimation. We use the following notations: denotes the logits for class , denotes the calibrated probability for class , means the dimension of the final output layer, and denotes the number of classes.
@@ 3.1 Density Softmax
Density Softmax Bui and Liu (2024a) obtains confidence by weighting logits with normalized log-likelihood from a trained density estimator. In this study, we use RealNVP Dinh et al. (2017) as the density estimator (details for the density estimator are in Appendix A). is the parameters of the density estimator; is the representation obtained from the input example ; is the normalized log-likelihood from the density estimator, then the corrected confidence is written as
In Density Softmax, the closer the normalized log-likelihood to zero, the closer the prediction to Uniform distribution Bui and Liu (2024a). Density Softmax achieves reasonable latency and competitive uncertainty performance with state-of-the-art methods at the cost of demanding the density estimator training and multiple base model training.
@@ 3.2 Density Aware Calibration (DAC)
DAC Tomani et al. (2023) scales the logits by using sample-dependent temperature
@@
are the weights for every layer of the base model, is the averaged distance from NN search on -th layer, and is the bias term. are optimized using the L-BFGS-B method Liu and Nocedal (1989) based on the loss in the validation set. In the original DAC paper, the uncertainty estimation performance tends to improve with the increase in the number of layer’s representation Tomani et al. (2023). Therefore, we report the performance when all the hidden representations in each layer of the base PLMs are used.
DAC avoids the problem of the high training cost in Density Softmax. However, both DAC and Density Softmax use only the density of the training data representations for uncertainty estimation. Considering that the uncertainty is mainly composed of epistemic uncertainty and aleatoric uncertainty, DAC and Density Softmax only represent the epistemic uncertainty because they use only the distance from neighbors or the data likelihood. Therefore, the uncertainty estimates could be further improved by introducing the concept of aleatoric uncertainty.
@@ 4 Proposed Method: -Nearest Neighbor Uncertainty Estimation (NN-UE)
Our proposed method, NN-UE, is based on the intuition that a model should give high confidence only when the predicted instance exists in the region with high density of training data and a high proportion of the model’s predicted label exists in the neighbors sharing the same label, as illustrated in Figure 1. To represent this intuition, we introduce a scoring method based on NN and labels of the neighbors in the following formula:
@@
is the number of neighbors from NN search, is the number of examples where the predicted label and the label of the neighbors are same, is the distance between the -th representation obtained by NN search and the representations of training data.222Note that NN-UE is also ”accuracy-preserving” same as DAC because is a scalar, not a class-wise score. The parameters , , and are optimized using the L-BFGS-B method based on the loss in the validation set. As with Density Softmax, the closer is to zero, the closer the prediction is to Uniform distribution, which allows us to better estimate confidence of the prediction. Unlike DAC, which relies only on the distances to neighbors and Density Softmax, which uses the log-likelihood from the density estimator, the label term in NN-UE increases the uncertainty when there is the high aleatoric uncertainty region that corresponds to the high variation of labels in neighbors of the predicted query representation.

### v2

@@ Abstract
Trustworthiness in model predictions is crucial for safety-critical applications in the real world. However, deep neural networks often suffer from the issues of uncertainty estimation, such as miscalibration. In this study, we propose -Nearest Neighbor Uncertainty Estimation (NN-UE), which is a new uncertainty estimation method that uses not only the distances from the neighbors, but also the ratio of labels in the neighbors. Experiments on sentiment analysis, natural language inference, and named entity recognition show that our proposed method outperforms the baselines and recent density-based methods in several calibration and uncertainty metrics. Moreover, our analyses indicate that approximate nearest neighbor search techniques reduce the inference overhead without significantly degrading the uncertainty estimation performance when they are appropriately combined.
@@ 1 Introduction
In order to deploy Deep Neural Networks (DNNs) including Pre-trained Language Models (PLMs) in safety-critical areas, uncertainty estimation (UE) is important. Improving the predictive uncertainty will calibrate the prediction Guo et al. (2017),111”Calibration” means the confidence of the prediction aligns with its accuracy. or enhance the selective prediction performance which reduces incorrect predictions by providing the option to abstain from the model prediction Galil et al. (2023). On the other hand, DNNs often fail to quantify the predictive uncertainty, for example, causing miscalibrated prediction Guo et al. (2017). Such UE performance problems can be mitigated by the PLMs, such as BERT Devlin et al. (2019) or DeBERTa He et al. (2021b), that are self-trained on vast amounts of data Ulmer et al. (2022); nevertheless, there remains considerable room for improvement Desai and Durrett (2020).
@@
To address the challenge of UE, multiple stochastic inferences such as MC Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017) are generally effective. On the other hand, these methods require multiple stochastic inferences for a single data instance, which leads to high computational cost, and makes them impractical for real world application. To balance reasonable predictive uncertainty with computational efficiency, Temperature Scaling Guo et al. (2017) , which scales logits by a temperature parameter, is commonly employed. Furthermore, density-based methods, such as Density Softmax Bui and Liu (2024) and Density Aware Calibration (DAC) Tomani et al. (2023), have demonstrated promising UE performance and inference costs by adjusting model outputs based on estimated density.
However, both Density Softmax and DAC only use the density of the training data. Relying on density alone can sometimes lead to overconfident predictions, even when such confidence is unwarranted. For instance, the neighbor of the input may contain many examples with labels that differ from the predicted label. In this situation, the prediction should obviously not be trusted. Therefore, we hypothesized that considering both the density and the label information of the neighbors will improve UE performance.
@@
In this study, we propose -Nearest Neighbor Uncertainty Estimation (NN-UE), a new density-based UE method that reflects nearest neighbor labels. As illustrated in Figure 1, NN-UE is designed to achieve the highest prediction confidence when the input and the nearest neighbors are both close in distance and share the same label as the predicted label. Our method weights logits according to the score from the distance between the input example and its neighbors in the datastore created by the training data and the ratio of the model’s predicted label matched with the labels in the neighbors. In addition, our method requires only a single forward inference of the model with almost no additional computational cost. The contributions of this research are as follows.
@@
First, our experiments show that NN-UE improves the UE performance of existing baselines in sentiment analysis, natural language inference, and named entity recognition in both in-domain and out-of-domain settings by combining neighbor label information and distances from neighbors. On the other hand, we also find that naive NN-UE makes less efficient for token-level tasks such as sequence-labeling based named entity recognition due to the execution of NN to each token.
Second, to mitigate the above latency problem in NN-UE, we show that approximate NN search or dimension reduction in NN-UE improves the inference speed without degrading UE performance much more, while combining them leads to degrading the UE performance.
Our code is available at https://github.com/wataruhashimoto52/knn_ue.
@@ 2 Related Work
Uncertainty Estimation for Natural Language Processing Tasks
Studies about UE for NLP tasks are limited when compared with those for image datasets. Kotelevskii et al. (2022) has shown excellent performance in classification with rejection tasks and out-of-distribution detection tasks using uncertainty scores using density estimation results. Vazhentsev et al. (2022) performed misclassification detection using Determinantal point processes Kulesza and Taskar (2012), spectral normalization, Mahalanobis distance and loss regularization in text classification and NER. However, these are still focusing only on the feature representation or the density, not the labels of the neighbors. He et al. (2024) proposed a framework that considers uncertainty between tokens in NER. However, the target task is limited to NER, and it is not for confidence calibration. Hashimoto et al. (2024) shows that simple data augmentation methods in NER can improve UE performance without additional inference costs, but its effectiveness is limited in the in-domain. Our NN-UE improves the UE performance in the in-domain and the out-of-domain classification and NER tasks using not only NN density but also neighbor labels.
@@
-Nearest Neighbor Language Model (NN-LM) Khandelwal et al. (2020) has been proposed, which performs linear interpolation of NN probability based on distance from neighbors and base model probability, in the language modeling task. -Nearest Neighbor Machine Translation (NN-MT) applied the NN-LM framework to machine translation Khandelwal et al. (2021). NN-LM and NN-MT have been successful because they enhance predictive performance through the memorization and use of rich token representations of pre-trained language models and mitigate problems such as a sparsity comes from low-frequency tokens Zhu et al. (2023). The main issue on NN-LM and NN-MT is the inference overhead, and there are several studies to solve this problem. He et al. (2021a) employs datastore compression, adaptive retrieval, and dimension reduction to reduce computational overhead with retaining perplexity. Deguchi et al. (2023) dramatically improves decoding speed by dynamically narrowing down the search area based on the source sentence. We investigate that whether UE performance in NN-UE can keep or not with reducing inference time by introducing some of the speed-up techniques established in kNN-LM/MT.
3 Preliminary
3.1 Definitions
In multiclass classification, we assume a dataset consisting of examples, where denotes its corresponding class label among possible classes.222In the case of sequence labeling, we can interpret the number of data as the product of the raw number of data instances and the sequence length. We use the trained neural network feature extractor and the classifier for classification, where . gives us the logits and we obtain the confidence .
@@ 3.2 Density Softmax
Density Softmax Bui and Liu (2024) obtains confidence by weighting logits with normalized log-likelihood from a trained density estimator. are the parameters of the density estimator; is the normalized log-likelihood from the density estimator, then the corrected confidence is written as
In Density Softmax, the closer the normalized log-likelihood to zero, the closer the prediction to Uniform distribution. Density Softmax achieves reasonable latency and competitive UE performance with state-of-the-art methods at the cost of demanding the density estimator training and multiple base model training.333Details for the density estimator in this study are in Appendix B.
@@ 3.3 Density Aware Calibration (DAC)
DAC is a confidence calibration method using multiple feature representations, which is similar to the NN-based out-of-distribution detection Sun et al. (2022). DAC Tomani et al. (2023) scales the logits by using sample-dependent temperature
@@
are the weights for every layer of the base model, is the averaged distance from NN search on -th layer, and is the bias term. are optimized using the L-BFGS-B method Liu and Nocedal (1989) based on the loss in the validation set. In the original DAC paper, the UE performance tends to improve with the increase in the number of layer representation Tomani et al. (2023). Therefore, we use all the hidden representations in each layer of the base PLMs.
DAC is a non-parametric method that makes not assumptions about the training data distribution unlike Density Softmax Bui and Liu (2024), which relies on some density estimators. On the other hand, the recent NN-based DAC still relies only on the distances to the neighbors. These methods do not take into account the label information of the input neighbors, which limits the improvement of the UE performance.
@@ 4 Proposed Method: -Nearest Neighbor Uncertainty Estimation (NN-UE)
The main idea of our proposed method, NN-UE, stems from the notion that the density-based UE methods can be further improved by using label information about the training data instances that make up the density.
In order to take into account the variance of neighbor labels, our NN-UE explicitly includes the label agreement information of the predicted instance and its neighbor examples when calculating the confidence. More specifically, we regard the prediction as more reliable only when the prediction is in a region where training data is dense and the predicted label and the labels of the data instances that make up the dense region are mostly the same, as illustrated in the right part of Figure 1. Otherwise, for example, if there is a lot of discrepancy in the neighbor labels and the predicted label, we treat the prediction as unreliable, indicated in the middle of Figure 1.
In our NN-UE, we introduce two terms: one related to the density of the training data and one related to the degree of agreement of the predicted data and neighbor labels. Confidence of -th label obtained by NN-UE is following the formula:
@@
is the number of neighbors from NN search, is the count when the predicted label and the label of the -th neighbor is same, is the distance between the -th representation obtained by NN search and the representations of training data.444Note that NN-UE is also ”accuracy-preserving” same as DAC because is a scalar, not a class-wise score. The parameters , , and are optimized using the L-BFGS-B method based on the loss in the validation set.
When the distance and label terms are smaller and is closer to zero, the closer the prediction is to Uniform distribution, which allows us to better estimate the confidence of the prediction. In this study, we also conduct experiments without the label term in Equation 5, to emphasize the importance of NN neighbor labels in UE. We summarize a diagram of NN-UE in Figure 2.